OpenStudy (anonymous):

which of the following are trigonometric indentities (check all that apply). a. 4cos(x)sin(x) = 2cos(x)+1-2sin(x) b. tan(x)cos(x)csc(x) = 1 c. (csc(x) - sin(x))/csc(x) = cos^2(x) d. 1 + cot(x) tan(y) = (sin(x+y))/sin(x)cos(y) Please help! I am really bad with identities!

OpenStudy (xapproachesinfinity):

b is an identity just check if the left had side equal to right hand side there is more than one

OpenStudy (xapproachesinfinity):

c as well is an identity

OpenStudy (anonymous):

but how would you figure it out

OpenStudy (anonymous):

because i would like to know how to prove and show an equation is an identity @xapproachesinfinity

OpenStudy (xapproachesinfinity):

use the identities you know like csc=1/sin tan=sin/cos ... you need to know the basic stuff

OpenStudy (xapproachesinfinity):

a is not good it is not an identity 4cosxsinx=2sin2x this cannot equal that left side stuff

OpenStudy (xapproachesinfinity):

of course you need to know that 2cosxsinx=sin2x

OpenStudy (xapproachesinfinity):

well now you have only one to work on recall sin(a+b)=sinacosb+sinbcosa

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

use that x=a y=b see if that satisfies the left hand side

OpenStudy (anonymous):

it doesn't

OpenStudy (xapproachesinfinity):

i agree

OpenStudy (xapproachesinfinity):

i think you are done:)

OpenStudy (anonymous):

so then the answers are b and c ?

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (xapproachesinfinity):

yes are you not paying attention hehe

OpenStudy (anonymous):

d is also an answer @xapproachesinfinity

OpenStudy (xapproachesinfinity):

No

OpenStudy (mathmath333):

one trick to check that is to put arbitrary values for x example \(\large\tt \begin{align} \color{black}{x=45^{\circ}\hspace{.33em}\\~\\}\end{align}\) and see if \(\bf LHS =RHS\)

OpenStudy (xapproachesinfinity):

i like this one :) never used in identities